OpenStudy (anonymous):

Solve for x. 12x - 8 = 8x - 20

OpenStudy (anonymous):

Subtract 8x from both sides, you'll get 4x - 8 = -20 Now add 8 to both sides 4x = -20 + 8 = -12 Divide both sides by 4 x = -12/4 = -3 :)

OpenStudy (jamesj):

First subtract 8x from each side 12x - 8x - 8 = 8x - 8x - 20 4x - 8 = -20 Now add 8 to both sides 4x -8 + 8 = -20 + 8 4x = -12 Then divide both sides by 4 4x/4 = -12/4 x = -3 At this stage you should now check this is a solution by substituting it back into the original equation. The left-hand side, LHS = 12x - 8 = 12(-3) - 8 = -44 And RHS = 8(-3) - 20 = -44. So LHS = RHS and the solution is correct.